You like to keep your investment risks at a 78–10–12 proportion (stocks-bonds-cash). After the first year, your $44,000 investment doubled in value to $88,000, with $70,400 in stock, $12,100 in bonds, and $5,500 in cash. How should you allocate your assets to maintain your original goals and rebalance your portfolio to retain the 78–10–12 proportion in investments? (Input all amounts as positive values.)
Answer #1

Stocks=88000*78%=68640.00
Bonds=88000*10%=8800.00
Cash=88000*12%=10560.00

Sell stocks worth 70400-68640=1760.00 and sell bonds worth 12100-8800=3300.00 and keep that amount in cash